OpenStudy (anonymous):

Find the functions a) f○g, b) g○f, c) f○f, and d)g○g and their domains 31.f(x)=x^2-1, g(x)=2x+1

OpenStudy (anonymous):

the composition of two polynomials is again a polynomial, so the answer to "their domains" is "all real numbers" for ever question

OpenStudy (anonymous):

How do you find the functions a-d

OpenStudy (anonymous):

a) f(g) = g^2-1=(2x+1)^2-1 b) g(f)=2f+1=2(x^2-1)+1

OpenStudy (anonymous):

Luckily I got that point but for a) they look like they simplified it further or in the back of the book they got 4x^2+4x and b)2x^2-1. Can you show me how they got those answers?

OpenStudy (anonymous):

(a+b)^2=a^2+2ab+b^2 (2x+1)^2=4x^2+4x+1 (2x+1)^2-1=4x^2+4x 2(x^2-1)+1=(2x^2-2)+1=2x^2-1

OpenStudy (anonymous):

Ok now I see thankyou. : )op